# An extrapolated and provably convergent algorithm for nonlinear matrix decomposition with the ReLU function

### v1

@@ 4.1 Previous works
Let us start with Latent-ReLU-NMD, since we use the algorithms designed to solve Latent-ReLU-NMD as baselines in the numerical comparison in Section 5. To our knowledge, all existing algorithms for Latent-ReLU-NMD are alternate optimization methods in which each block of variables is updated sequentially [16, 17, 32, 31, 40] and lack convergence guarantees.
The first approach is the so-called Naive-ReLU-NMD, introduced by Saul in [34], which employs a basic alternating scheme over $Z$ and $\Theta$ with closed-form solutions for both. Similar approaches have been studied in matrix completion, see for example [5, 8, 10, 15]. At each iteration, firstly $\Theta$ is fixed and then the optimal $Z$ is computed solving the subproblem
@@
$Z_{ij}=\begin{cases}X_{ij}\quad&\mbox{if }X_{ij}>0\\ \min(0,\Theta_{ij})&\mbox{if }X_{ij}=0.\\ \end{cases}$
Once $Z$ is updated, $\Theta$ is updated solving
$\operatorname{argmin}_{\Theta}\lVert Z-\Theta\rVert_{F}^{2}\quad\mbox{such that}\quad\operatorname{rank}(\Theta)\leq r,$
whose optimal solution(s) can be obtained with the TSVD of $Z$. Despite its simplicity, this scheme is highly reliable in practice, with the most costly operation at each iteration being the computation of a TSVD.
An accelerated, heuristic version of the Naive-NMD scheme was proposed in [35]. Specifically, the so-called Aggressive-ReLU-NMD (A-NMD) employs an extrapolation step for both $Z$ and $\Theta$ while the extrapolation parameter changes adaptively at each iteration, depending on the value of the residual. A-NMD is an example where, despite its lack of theoretical guarantees, extrapolation can be considerably faster [2, 43].
Another algorithm for solving Latent-ReLU-NMD is the so-called expectation-minimization NMD (EM-NMD) [34]. This approach is based on a Gaussian latent variable model parametrized by the low-rank matrix $\Theta$ and a parameter $\sigma$. For each entry of the original matrix, a Gaussian latent variable $Z_{ij}=\mathcal{N}(\Theta_{ij},\sigma^{2})$ is sampled. The method estimates $\Theta$ and $\sigma$ by maximizing the likelihood of the data with respect to $\Theta$ and $\sigma$. The optimization is carried out using the Expectation-Minimization (EM-NMD) scheme that consists of two main steps: the E-step which computes posterior mean and variance of the Gaussian latent variable model, and the M-step re-estimates the parameters of the model from the posterior mean and variance.
The BCD scheme for solving 3B-ReLU-NMD in (5), referred to as BCD-NMD, consists of an alternating optimization procedure over three blocks of variables $(Z,W,H)$. BCD-NMD belongs to the class of exact BCD methods, where a global optimizer for each of the subproblem is computed sequentially. Let $A^{\dagger}$ denote the Moore-Penrose inverse of $A$ and let $(H^{k},Z^{k},W^{k})$ be the current iteration. The $(k+1)$-th step of BCD-NMD is as follows:
@@
$H^{k+1}=$ $\operatorname{argmin}_{H}\lVert Z^{k+1}-W^{k+1}H\rVert_{F}^{2}=((W^{k+1})^{T})^{\dagger}Z^{k+1}.$
The solution of the subproblems for $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ are analogous. In particular, the first is separable by rows of $W$, leading to $m$ least squares problems in $r$ variables with the data vector of dimension $n$, while the subproblem for $H$ is separable by column and equivalent to $n$ least squares problems of $r$ variables and data vector of dimension $m$. Both these subproblems do not necessarily have a unique solution. In fact, uniqueness is guaranteed only if $W^{k}$ and $H^{k}$ remain full rank throughout the iterations. Alternatively, uniqueness is ensured by adding a Tikhonov regularization term or a proximal term in each of the subproblems [16, 17, 3]; however, it is not necessary to prove the convergence of the BCD-NMD algorithm.
@@
The overall BCD-NMD algorithm is reported in Algorithm 1 and has a computational cost of $O(r|\Omega^{C}|+(m+n)r^{2})$.
Algorithm 1 BCD-NMD: block coordinate descent for ReLU-NMD
@@
Similarly to the Naive scheme for Latent-ReLU-NMD, adding extrapolation might be beneficial also for BCD-NMD. In particular, the work [35] proposes to perform an extrapolation step on the latent variable $Z$ and on the low-rank product $WH$, resulting in the so-called 3B-NMD algorithm. Moreover, Wang et al. [41, 42] added a Tikhonov regularization for both the variables $W$ and $H$ in the 3B-ReLU-NMD formulation. Their algorithm is dubbed as ReLU-NMD-T and it is based on the 3B-NMD scheme but performs two more extrapolation steps following the computation of $W$ and $H$. Even though numerical experience shows that both the accelerated algorithms outperform the original BCD-NMD, convergence still remains an open problem. For the numerical experiments in Section 5, we consider only 3B-NMD since we are interested in comparing algorithms that solve the non-regularized 3B-ReLU-NMD.
4.2 Convergence of BCD-NMD
In this section, we discuss the convergence to a stationary point of BCD-NMD (Algorithm 1) which, to the best of our knowledge, has not been studied before. Indeed, our goal is to prove that BCD-NMD is convergent under the assumption that a limit point of the sequence generated by the scheme exists. We proceed by showing that the method falls into the more general framework of exact block coordinate descent (BCD), where one of the subproblems is strictly quasi-convex.
Specifically, the general BCD scheme is studied by Grippo et al. [16] when the objective function is continuously differentiable over the Cartesian product of $m$ closed, convex sets, one for each block of variables. We first observe that the 3B-ReLU-NMD objective function is continuously differentiable and defined over three blocks of variables $Z$, $W$, and $H$. The latent variable $Z$ is the only constrained variable, and it is easy to show that the constraint set in (17) is convex. Therefore, the 3B-ReLU-NMD is a particular instance of the more general problem considered in [16].
@@ 4.2 Convergence of BCD-NMD
Any limit point of the sequence generated by BCD-NMD (Algorithm 1) is a stationary point of problem (5).
@@
The global convergence of the BCD scheme for a general function with $m$ blocks of variables is analyzed under generalized convexity assumptions in [16] (Proposition 5). The key hypotheses are: 1) the existence of a limit point of the sequence generated by the BCD algorithm, and 2) strict quasi-convexity of at least $m-2$ subproblems. Therefore, since 3B-ReLU-NMD has three blocks of variables, we just need one subproblem to be strictly quasi-convex. The first subproblem in the variable $Z$ is a convex, quadratic problem over a convex set, making it strictly quasi-convex, implying the convergence of BCD-NMD.
The assumption of the existence of the limit point is an undesirable hypothesis. However, it is necessary because the level sets of the objective function are not bounded. Therefore, similarly to Latent-ReLU-NMD, 3B-ReLU-NMD might have instances in which the minimum is not attained (see Theorem 3.3). This means that, in pathological cases, the sequence generated by BCD-NMD may diverge simply because the infimum of the function is not attained by any finite value. Thus, the hypothesis regarding the existence of a limit point of the BCD-NMD sequence cannot be weakened. One possible way to overcome this limitation is by adding a regularization term to the objective function (5), which makes the level sets compact, as in [41]. However, this is not the purpose of this work which focuses on the non-regularized formulation.
4.3 Extrapolated block coordinate descent NMD (eBCD-NMD)
We present now a new extrapolated version of Algorithm 1, namely extrapolated BCD-NMD (eBCD-NMD). Our goal is to accelerate BCD-NMD while preserving convergence to a stationary point under mild assumptions. The eBCD-NMD method is an adaptation of the well-known LMaFit algorithm [43] for matrix completion. The standard matrix completion formulation in [43] is similar to the one we consider in (5), although it does not include the inequality constraint $P_{\Omega^{C}}(Z)\leq 0$, which models the crucial information that the missing entries are negative. Consequently, we derive a new optimization scheme, differing from the one in [43], as the update of $Z$ presents an additional nonlinear term; see equation (18). Moreover, we relax the hypothesis in [43] that require $\{W^{k}\}$ and $\{H^{k}\}$ to remain full-rank for all iterations. We present two different versions of the eBCD-NMD algorithm that, starting from the same point, produce the same low-rank approximation $W^{k}H^{k}$, yet the factors $W^{k}$ and $H^{k}$ vary. In what follows, we omit the iteration index, $k$, to simplify the notation, as we focus on a single iteration eBCD-NMD. We assume we have a point $(Z,W,H)$ from which we perform one iteration.
4.3.1 First version of eBCD-NMD
Let $\alpha\geq 1$, then the first version of the eBCD-NMD scheme uses the following update: given the iterate $(Z,W,H)$, it computes
@@ 4.3.1 First version of eBCD-NMD
where $\left(\widehat{Z}(\alpha),\widehat{W}(\alpha),\widehat{H}(\alpha)\right)$ provides the next iterate. Similarly to BCD-NMD, the main computational cost for the scheme in (18) is solving the two least squares problems for computing $\widehat{W}(\alpha)$ and $\widehat{H}(\alpha)$. The optimization scheme in (18) differs from the one in (16) because of the new variable $Z_{\alpha}$ which is used in place of the original latent variable $Z$ to update $W$ and $H$. Moreover, if we assume the matrix $H$ to be full-rank, we can derive an intuitive interpretation of the update of $Z_{\alpha}$ as an indirect extrapolation step. Since $H$ is full-rank we have $H^{\dagger}=H^{T}(HH^{T})^{-1}$ and thus
@@
where $W_{\text{BCD}}$ is the update of the variable $W$ that we would have had if we performed one step of the BCD-NMD scheme (16). Moreover, if we define $\beta=\alpha-1$, then
(19) $\widehat{W}(\alpha)=\alpha W_{\text{BCD}}+(1-\alpha)W=(1+\beta)W_{\text{BCD}}-\beta W=W_{\text{BCD}}+\beta(W_{\text{BCD}}-W),$
that is, an extrapolation step for the BCD-NMD update of $W$, provided that $\alpha\in(1,2)$, that means $\beta\in(0,1)$. Unfortunately, such an intuitive interpretation does not apply to the update of the variable $H$. Note that the step in (19) is properly referred to as an extrapolation step if $\beta$ is in the interval (0,1) or equivalently $\alpha$ is in the interval (1,2), but our scheme potentially allows for larger values that exceed the upper bound.
4.3.2 Improved version of eBCD-NMD
The second version of the eBCD-NMD does not require solving any least squares problem and the main computational cost is associated with a single QR factorization of a matrix in $\mathbb{R}^{m\times r}$. We specify that this scheme can be adapted to design alternative versions of BCD-NMD and 3B-NMD as well. We state the following lemma from [43] that allows us to derive the second version of eBCD-NMD.
@@ 4.3.2 Improved version of eBCD-NMD
Let $UDV^{T}=H$ be the economy form SVD of $H$, then
$\widehat{W}(\alpha)=Z_{\alpha}H^{\dagger}=Z_{\alpha}VD^{\dagger}U^{T}.$
Moreover, since $Z_{\alpha}H^{T}=Z_{\alpha}VDU^{T}$, we get $\mathcal{R}(\widehat{W}(\alpha))=\mathcal{R}(Z_{\alpha}H^{T})$.
Lemma 4.3 shows that the range space of $\widehat{W}(\alpha)$ can be obtained from that of $Z_{\alpha}H^{T}$. Let $Q(\alpha)$ be a matrix whose columns contain an orthogonal basis for $\mathcal{R}(Z_{\alpha}H^{T}).$ Then, using Lemma 4.3 and from the second and third lines in (18), we get
@@
This shows that the low-rank approximation produced by the eBCD-NMD is nothing else than the orthogonal projection of $Z_{\alpha}$ into $\mathcal{R}(Z_{\alpha}H^{T})$. Justified by the relation in (20), we now state the improved version of eBCD-NMD. Given an iterate $(Z,W,H)$, we compute
@@
$Q(\alpha)\rightarrow\text{ Orthogonal basis of }\mathcal{R}(Z_{\alpha}H^{T}),$
$W(\alpha)=Q(\alpha),$
@@
to obtain the next iterate $\left(Z(\alpha),W(\alpha),H(\alpha)\right)$. Equation (20) guarantees that the low-rank approximations computed by the scheme in (18) and in (21) are equal, that is $\widehat{W}(\alpha)\widehat{H}(\alpha)=W(\alpha)H(\alpha)$. In practice, $Q(\alpha)$ can be computed from the economy QR decomposition of $Z_{\alpha}H^{T}$. Note that if $\operatorname{rank}(Z_{\alpha}H^{T})=r_{1}<r$, the first $r_{1}$ columns of the $Q$ matrix from the economy QR decomposition of $Z_{\alpha}H^{T}$ are not, in general, an orthogonal basis of the range space of the matrix. To ensure this property, a modified version of the QR decomposition using column pivoting must be used instead. Specifically, there exists $Q(\alpha)\in\mathbb{R}^{m\times r_{1}}$, containing an orthogonal basis of $Z_{\alpha}H^{T}$, such that $Z_{\alpha}H^{T}\Pi=Q(\alpha)R(\alpha)$, where $\Pi$ is a permutation matrix and $R(\alpha)\in\mathbb{R}^{r_{1}\times r}$ [14]. Therefore, we potentially allow for a drop in the rank of the approximation whenever $Z_{\alpha}H^{T}$ is rank-deficient. However, we have never encountered this case in practical situations because $Z_{\alpha}H^{T}$ is likely to have rank at least $r$ while the factors $Q(\alpha)$ and $H(\alpha)$ have rank exactly $r$.
@@
We need now to determine a rule to select the extrapolation parameter $\alpha$ at each iteration. Extrapolating a variable means adding information from the past iterate, thus the extrapolation parameter $\alpha$ affects how much relevant is the past information in the current update. Of course, not all the values of $\alpha$ are acceptable to progress in minimizing the residual, and a good choice of the parameter is crucial. We present now the updating scheme of the extrapolation parameter $\alpha$, inspired by the one introduced in [43]. The main idea is to choose the parameter $\alpha$ such that the residual of the function is decreased sufficiently between two consecutive iterations.
@@
If $\delta(\alpha)<1$, we declare our update successful and the new iterate is set to $(W^{k+1},H^{k+1},Z^{k+1})=(W(\alpha),H(\alpha),Z(\alpha))$. Furthermore, let $\bar{\delta}$ be a fixed threshold, if $\delta(\alpha)>\bar{\delta}\in(0,1)$, meaning we do not consider the decrease of the residual fast enough, we increase $\alpha$ by setting $\alpha\leftarrow\min(\alpha+\mu,\bar{\alpha})$, where $\mu=\max(\mu,0.25(\alpha-1))$. If $\alpha$ reaches the upper bound $\bar{\alpha}$, we restart the parameter by setting $\alpha=1$. This correction does not appear in [43], but we added it to guarantee the convergence of the our algorithm. On the contrary, if $\delta(\alpha)\geq 1$, the step is declared unsuccessful, we do not accept the update, and we compute a standard BDC-NMD update instead by setting $\alpha=1$. Consequently, either we find an $\alpha\in(1,\bar{\alpha})$ that allows a decrease of the residual or we recover one iteration of the BCD-NMD, which is guaranteed to decrease the residual. Therefore, the new iterate of the algorithm produces a lower value in the residual with respect to the previous iteration.
The overall eBCD-NMD scheme is described in Algorithm 2, where we consider the updating scheme in (21). We choose the improved version of eBCD-NMD because (a) it exhibits faster performance in practice; (b) computing one QR factorization is in general more stable than solving two consecutive least squares problems; (c) one of the factors in the decomposition is orthogonal and this helps demonstrate crucial properties in the convergence analysis, such as boundness of the sequence generated by the eBCD-NMD algorithm. Let us also mention that, in [43], convergence is proved for a scheme similar to the one in (18), instead of the scheme in (21) as done in this paper.
Algorithm 2 eBCD-NMD
0:  $X$, $Z^{0}$,$W^{0}$, $H^{0}$, $\bar{\alpha}$ (default $=4$), $\mu$ (default $=0.3$), $\bar{\delta}$ (default $=0.8$), maxit
@@
4:     Compute $Q(\alpha_{k})$ orthogonal basis of $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$

### v2

@@ 4.1 Previous works
Let us start with the algorithm designed for Latent-RMD since we use them as baselines in the numerical comparison in Section 5. To our knowledge, all existing algorithms are alternating optimization methods in which each block of variables is updated sequentially [grippo2000convergence, grippo2023introduction, patriksson1998decomposition, powell1973search, tseng1991decomposition] and lack convergence guarantees.
The first approach is the so-called Naive algorithm, introduced by Saul in [saul2022nonlinear], which employs a basic alternating scheme over $Z$ and $\Theta$ with closed-form solutions for both. Similar approaches have been studied in matrix completion, see for example [balzano2010online, boumal2015low, dai2012geometric, goyens2023nonlinear]. At each iteration, firstly $\Theta$ is fixed and then the optimal $Z$ is computed solving the subproblem
@@
[…]
Once $Z$ is updated, $\Theta$ is updated solving $\operatorname{argmin}_{\Theta}\lVert Z-\Theta\rVert_{F}^{2}$ such that $\operatorname{rank}(\Theta)\leq r$, whose optimal solution(s) can be obtained with the TSVD of $Z$. Despite its simplicity, this scheme is highly reliable in practice, with the most costly operation at each iteration being the computation of a TSVD.
An accelerated, heuristic version of the Naive scheme was proposed in [seraghiti2023accelerated]. Specifically, the so-called Aggressive Naive (A-Naive) algorithm, employs an extrapolation step for both $Z$ and $\Theta$ while the extrapolation parameter changes adaptively at each iteration, depending on the value of the residual. A-Naive is an example where, despite its lack of convexity and theoretical guarantees, extrapolation can be considerably faster the non accelerated counterpart [ang2019accelerating, wen2012solving].
Another algorithm for solving Latent-RMD is the expectation-minimization (EM) algorithm [saul2022nonlinear]. This approach is based on a Gaussian latent variable model parametrized by the low-rank matrix $\Theta$ and a parameter $\sigma$. For each entry of the original matrix, a Gaussian latent variable $Z_{ij}=\mathcal{N}(\Theta_{ij},\sigma^{2})$ is sampled. The method estimates $\Theta$ and $\sigma$ by maximizing the likelihood of the data with respect to $\Theta$ and $\sigma$. The optimization is carried out using the expectation-minimization scheme that consists of two main steps: the E-step which computes posterior mean and variance of the Gaussian latent variable model, and the M-step re-estimates the parameters of the model from the posterior mean and variance.
Finally, we present the BCD algorithm for solving the 3B-RMD model. In what follows, we use the formulation of 3B-RMD introduced in (4) rather than the equivalent one in (3). The BCD algorithm consists of an alternating optimization procedure over three blocks of variables $(Z,W,H)$, and computes a global optimizer for each of the subproblems sequentially. Let $A^{\dagger}$ denote the Moore-Penrose inverse of $A$, and let $(H^{k},Z^{k},W^{k})$ be the current iteration. The $(k+1)$-th step of BCD is as follows:
@@
$H^{k+1}=$ $\operatorname{argmin}_{H}\lVert Z^{k+1}-W^{k+1}H\rVert_{F}^{2}=((W^{k+1})^{T})^{\dagger}Z^{k+1},$
where $P_{\Omega}$ and $P_{\Omega}^{C}$ are defined as in (4). The solutions of the subproblems for $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ are analogous. In particular, the first is separable by rows of $W$, leading to $m$ least squares problems in $r$ variables with data vector of dimension $n$, while the subproblem for $H$ is separable by column, and equivalent to $n$ least squares problems of $r$ variables with data vector of dimension $m$. Both these subproblems do not necessarily have unique solutions. In fact, uniqueness is guaranteed only if $W^{k}$ and $H^{k}$ remain full rank throughout the iterations. Alternatively, uniqueness is ensured by adding a Tikhonov regularization term or a proximal term in each of the subproblems [attouch2010proximal, grippo2000convergence, grippo2023introduction]. However, regularization might slow down the algorithm, and it is not necessary to prove the convergence of the BCD scheme in this context.
@@
The overall BCD algorithm is reported in Algorithm 1 and has a computational cost of $O(r|\Omega^{C}|+(m+n)r^{2})$.
Algorithm 1 BCD: block coordinate descent for 3B-RMD
@@
Similarly to the Naive scheme for Latent-RMD, adding extrapolation might be beneficial also for BCD. In particular, the work [seraghiti2023accelerated] proposes to perform an extrapolation step on the latent variable $Z$ and on the low-rank product $WH$, resulting in the so-called extrapolated 3B (e3B) algorithm. Moreover, Wang et al. [wang2024momentum, wang2025accelerated] added a Tikhonov regularization for both the variables $W$ and $H$. Their algorithm is based on the e3B scheme but performs two additional extrapolation steps following the computation of $W$ and $H$. Even though numerical experience shows that both accelerated algorithms outperform the original BCD, convergence still remains an open problem.
4.2 Convergence of BCD
In this section, we discuss the convergence to a stationary point of the BCD algorithm (Algorithm 1), which, to the best of our knowledge, has not been studied before in the context of RMD. Indeed, our goal is to prove that BCD is convergent under the assumption that a limit point of the sequence generated by the scheme exists. We proceed by showing that the method falls into the more general framework of exact BCD, where one of the subproblems is strictly quasi-convex. Specifically, the general BCD scheme is studied by Grippo and Sciandrone [grippo2000convergence] when the objective function is continuously differentiable over the Cartesian product of $m$ closed, convex sets, one for each block of variables. We first observe that the 3B-RMD objective function is continuously differentiable and defined over three blocks of variables $Z$, $W$, and $H$. The latent variable $Z$ is the only constrained variable, and it is easy to show that the constraint set in (16) is convex. Therefore, 3B-RMD is a particular instance of the more general problem considered in [grippo2000convergence].
@@ 4.2 Convergence of BCD
Any limit point of the sequence generated by BCD (Algorithm 1) is a stationary point of problem (4).
@@
The global convergence of the BCD scheme for a general function with $m$ blocks of variables is analyzed under generalized convexity assumptions in [grippo2000convergence] (Proposition 5). The key hypotheses are: 1) the existence of a limit point of the sequence generated by the BCD algorithm, and 2) strict quasi-convexity of at least m2m-2 subproblems. Therefore, since 3B-RMD has three blocks of variables, we just need one subproblem to be strictly quasi-convex. The first subproblem in the variable $Z$ is a convex, quadratic problem over a convex set, making it strictly quasi-convex, implying the convergence to a stationary point.
The assumption of the existence of the limit point is an undesirable hypothesis. However, it is necessary because the level sets of the objective function are not bounded. Therefore, similarly to Latent-RMD, 3B-RMD might have instances in which the minimum is not attained (see Theorem 3.3). This means that, in pathological cases, the sequence generated by BCD may diverge simply because the infimum of the function is not attained by any finite value. Thus, the hypothesis regarding the existence of a limit point of the BCD sequence cannot be weakened. One possible way to overcome this limitation is by adding a regularization term to the objective function (4), which makes the level sets compact, as in [wang2024momentum]. However, this is not the purpose of this work which focuses on the non-regularized formulation.
4.3 Extrapolated BCD (eBCD)
We want to accelerate the BCD algorithm while preserving convergence to a stationary point under mild assumptions. The eBCD method is an adaptation of the well-known LMaFit algorithm [wen2012solving] for matrix completion. The standard matrix completion formulation in [wen2012solving] is similar to the one we consider in (4), although it does not include the inequality constraint $P_{\Omega^{C}}(Z)\leq 0$, which models the crucial information that the missing entries are negative. Consequently, we derive a new optimization scheme, differing from the one in [wen2012solving], as the update of $Z$ presents an additional nonlinear term; see equation (17). Moreover, we relax the hypothesis in [wen2012solving] that requires $\{W^{k}\}$ and $\{H^{k}\}$ to remain full-rank for all iterations. We present two different versions of the eBCD algorithm that, starting from the same point, produce the same low-rank approximation $W^{k}H^{k}$, yet the factors $W^{k}$ and $H^{k}$ vary. In what follows, we omit the iteration index, $k$, to simplify the notation, as we focus on a single iteration of the eBCD algorithm. We assume we have a point $(Z,W,H)$ from which we perform one iteration.
4.3.1 First version of eBCD
Let $\alpha\geq 1$, then the first version of the eBCD scheme uses the following update: given the iterate $(Z,W,H)$, it computes
@@ 4.3.1 First version of eBCD
where $\left(\widehat{Z}(\alpha),\widehat{W}(\alpha),\widehat{H}(\alpha)\right)$ provides the next iterate. Similarly to BCD, the main computational cost for the scheme in (17) is solving the two least squares problems for computing $\widehat{W}(\alpha)$ and $\widehat{H}(\alpha)$. The optimization scheme in (17) differs from the one in (15) because of the new variable $Z_{\alpha}$ which is used in place of the original latent variable $Z$ to update $W$ and $H$. Moreover, if we assume the matrix $H$ to be full-rank, we can derive an intuitive interpretation of the update of $Z_{\alpha}$ as an indirect extrapolation step. Since $H$ is full-rank we have $H^{\dagger}=H^{T}(HH^{T})^{-1}$ and thus
@@
where $W_{\text{BCD}}$ is the update of the variable $W$ that we would have had if we performed one step of the BCD scheme (15). Moreover, if we define $\beta=\alpha-1$, then
(18) $\widehat{W}(\alpha)=\alpha W_{\text{BCD}}+(1-\alpha)W=(1+\beta)W_{\text{BCD}}-\beta W=W_{\text{BCD}}+\beta(W_{\text{BCD}}-W).$
Thus, line (18) represents an extrapolation step for $W_{\text{BCD}}$, provided that $\alpha\in(1,2)$, that means $\beta\in(0,1)$. Unfortunately, such an intuitive interpretation does not apply to the update of the variable $H$. Note that the step in (18) is properly referred to as an extrapolation step if $\beta$ is in the interval (0,1) or equivalently, $\alpha$ is in the interval (1,2), but our scheme potentially allows for larger values that exceed the upper bound.
4.3.2 Improved version of eBCD
The second version of the eBCD scheme does not require solving any least squares problem and the main computational cost is associated with a single QR factorization of a matrix in $\mathbb{R}^{m\times r}$. We specify that this scheme can be adapted to design alternative versions of BCD and e3B as well. We state the following lemma from [wen2012solving] that is used in what follows.
@@ 4.3.2 Improved version of eBCD
Let $UDV^{T}=H$ be the economy form SVD of $H$, then $\widehat{W}(\alpha)=Z_{\alpha}H^{\dagger}=Z_{\alpha}VD^{\dagger}U^{T}$. Moreover, since $Z_{\alpha}H^{T}=Z_{\alpha}VDU^{T}$, we get $\mathcal{R}(\widehat{W}(\alpha))=\mathcal{R}(Z_{\alpha}H^{T})$.
Lemma 4.3 shows that the range space of $\widehat{W}(\alpha)$ can be obtained from that of $Z_{\alpha}H^{T}$. Let $Q(\alpha)$ be a matrix whose columns form an orthonormal basis for $\mathcal{R}(Z_{\alpha}H^{T}).$ Then, using Lemma 4.3 and from the second and third lines in (17), we get
@@
This shows that the low-rank approximation produced by the eBCD is nothing else than the orthogonal projection of $Z_{\alpha}$ into $\mathcal{R}(Z_{\alpha}H^{T})$. Justified by the relation in (19), we now state the improved version of eBCD. Given an iterate $(Z,W,H)$, we compute
@@
$W(\alpha)=Q(\alpha){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ \leftarrow}}\text{ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0} Orthonormal}} basis of }\mathcal{R}(Z_{\alpha}H^{T}),$
@@
to obtain the next iterate $\left(Z(\alpha),W(\alpha),H(\alpha)\right)$. Equation (19) guarantees that the low-rank approximations computed by the scheme in (17) and in (20) are equal, that is $\widehat{W}(\alpha)\widehat{H}(\alpha)=W(\alpha)H(\alpha)$. In practice, $Q(\alpha)$ can be computed from the economy QR decomposition of $Z_{\alpha}H^{T}$. Note that if $\operatorname{rank}(Z_{\alpha}H^{T})=r_{1}<r$, the first $r_{1}$ columns of the $Q$ matrix from the economy QR decomposition of $Z_{\alpha}H^{T}$ are not, in general, an orthonormal basis of the range space of the matrix. To ensure this property, a modified version of the QR decomposition using column pivoting must be used instead. Specifically, there exists $Q(\alpha)\in\mathbb{R}^{m\times r_{1}}$, containing an orthonormal basis of $Z_{\alpha}H^{T}$, such that $Z_{\alpha}H^{T}\Pi=Q(\alpha)R(\alpha)$, where $\Pi$ is a permutation matrix and $R(\alpha)\in\mathbb{R}^{r_{1}\times r}$ [golub2013matrix]. Therefore, we potentially allow for a drop in the rank of the approximation whenever $Z_{\alpha}H^{T}$ is rank-deficient. However, we have never encountered this case in practical situations because $Z_{\alpha}H^{T}$ is likely to have rank at least $r$ while the factors $Q(\alpha)$ and $H(\alpha)$ have rank exactly $r$.
@@
We need to determine a rule to select the extrapolation parameter $\alpha$ at each iteration. Extrapolating a variable means adding information from the past iterate; thus, the extrapolation parameter $\alpha$ affects the relevance of the past information in the current update. Of course, not all the values of $\alpha$ are acceptable to progress in minimizing the residual, and a good choice of the parameter is crucial. The updating scheme of the extrapolation parameter $\alpha$, is inspired by the one introduced in [wen2012solving]. The main idea is to choose the parameter $\alpha$ such that the residual of the function is sufficiently decreased between two consecutive iterations.
@@
If $\delta(\alpha)<1$, we declare our update successful, and the new iterate is set to $(W^{k+1},H^{k+1},Z^{k+1})=(W(\alpha),H(\alpha),Z(\alpha))$. Furthermore, let $\bar{\delta}$ be a fixed threshold, if $\delta(\alpha)>\bar{\delta}\in(0,1)$, meaning we do not consider the decrease of the residual fast enough, we increase $\alpha$ by setting $\alpha\leftarrow\min(\alpha+\mu,\bar{\alpha})$, where $\mu=\max(\mu,0.25(\alpha-1))$. If $\alpha$ reaches the upper bound $\bar{\alpha}$, we restart the parameter by setting $\alpha=1$. This correction does not appear in [wen2012solving], but we added it to guarantee the convergence of our algorithm. On the contrary, if $\delta(\alpha)\geq 1$, the step is declared unsuccessful, we do not accept the update, and we compute a standard BCD update instead by setting $\alpha=1$. Consequently, either we find an $\alpha\in(1,\bar{\alpha})$ that allows a decrease of the residual, or we recover one iteration of the BCD, which is guaranteed to decrease the residual. Therefore, the new iterate of the algorithm produces a lower value in the residual with respect to the previous iteration.
The overall eBCD scheme is described in Algorithm 2, where we consider the updating scheme in (20). We choose the improved version of eBCD because (a) it exhibits faster performance in practice; (b) computing one QR factorization is, in general, more stable than solving two consecutive least squares problems; (c) one of the factors in the decomposition is orthogonal and this helps demonstrate crucial properties in the convergence analysis, such as boundness of the sequence generated by the eBCD algorithm. Let us also mention that, in [wen2012solving], convergence is proved for a scheme similar to the one in (17), instead of the scheme in (20) as done in this paper.
Algorithm 2 eBCD for 3B-RMD
0:$X$, $Z^{0}$$W^{0}$, $H^{0}$, $\bar{\alpha}$ (default =4=4), $\mu$ (default =0.3=0.3), $\bar{\delta}$ (default =0.8=0.8), maxit
@@
4:  Compute $Q(\alpha_{k})$ orthonormal basis of $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$